# Probing quantum properties of black holes with a Floquet-driven optical lattice simulator

### v1

@@ Abstract
To demonstrate this, we plot in Fig. 3(c) the time evolution of the population density $n_{h}=\langle c_{l_{h}}^{\dagger}c_{l_{h}}\rangle$ at the event horizon site $l_{h}$. The density $n_{h}$ is initially zero until the wave packet has reached the horizon, after which it reaches a maximum at a time $t_{f}$ before decreasing again. In Fig. 3(d) we plot $t_{f}$ against the temperature $\tilde{T}_{H}$ extracted from the thermalised spectrum and find the simple inverse relationship
@@
with a proportionality constant $\gamma$, directly linking Hawking temperature and time dynamics. This effect also transpires in the inset of Fig. 3(c), where the density evolutions at the event horizon are collapsed by a linear rescaling of the time axis. Note that $\gamma$ depends on the particular experimental conditions, in particular the initial position of the walker. We highlight that monitoring density at a single site is sufficient since the Hawking temperature is controlled by the local curvature at the horizon [3, 4, 6, 5, 74]. Alternatively, one can measure the total probability density outside the horizon, i.e. $P_{\textrm{out}}=\sum_{l=l_{h}+1}^{l=L}n_{l}$, where $n_{l}=\langle c_{l}^{\dagger}c_{l}\rangle$, which provides a direct measure for the total intensity of the Hawking radiation. Within the linear profile of Eq. (3), scaling the encoded Hawking temperature $T_{H}$ amounts to an equivalent global scaling of the tunnelling coefficients $\kappa_{l}$, which is directly equivalently to an inverse scaling of time and hence explains the behaviour observed in (8).
@@
Our simulator is described by a single-particle Hamiltonian, hence we do not anticipate thermalisation or scrambling. However, due to the specific tunnelling profile that simulates the black hole horizon, components of quantum states that tunnel through the horizon appear thermal for the degrees of freedom outside the black hole even for a pure initial quantum state. This corresponds to the notion that the Hawking radiation loses memory of the initial state of the black hole when observed from the outside. This is manifested in our simulator by obtaining a thermal spectrum outside the black hole with the same Hawking temperature regardless of the initial position of the walker inside the black hole.
@@
Figure 4: (a) Error resilience to random on-site potentials (blue) with $\varepsilon=0.1$, additive random nearest-neighbour (nn) tunnelling error (orange) ($\varepsilon=0.1$) and random next-nearest neighbour (nnn) tunnelling with $\varepsilon=0.01$ (green). These results are averaged over time, initialised sites and 100 realizations for a system with size $L=100$ and horizon $l_{h}=50$. The error bar indicates the standard deviation across different realizations. The theoretical Hawking temperature value is represented by the black lines. (b) Density-density interaction of strength $V$ for a system size of $L=30$ and $l_{h}=15$.
@@
Resilience to errors:– We now investigate the sensitivity of the extracted Hawking temperature to errors pertinent to Floquet-driven optical lattice experiments. Specifically, we extract the Hawking temperature $\tilde{T}_{H}$ as per the method used in Fig. 3(b), after introducing noise to the system Hamiltonian in Eq. 1. We model random on-site potentials as an additional term of the form $J_{0}\sum_{l=1}^{N}\varepsilon_{l}n_{l}$ in the Hamiltonian with the coefficients $\varepsilon_{l}$ drawn randomly from a standard normal distribution with mean zero and variance $\varepsilon$. Despite the existence of random potentials, the essential features of the extracted $T_{H}$ persist even with an impurity strength of $\varepsilon=0.1$ for a wide range of $\alpha$’s, as shown in Fig. 4(a). We next evaluate nearest-neighbour ($k=1$), and next-nearest neighbour ($k=2$) tunnelling errors, represented by $J_{0}\sum_{l=1}^{N}\varepsilon_{l}(\hat{c}_{l}^{\dagger}\hat{c}_{l+k}+\hat{c}_{l+k}^{\dagger}\hat{c}_{l})$. These errors can reflect higher-order corrections beyond Eq. (6), see [51]. Fig. 4(a) shows that the extracted Hawking temperature remains stable for error strength $\varepsilon=0.1$ for $k=1$ and $\varepsilon=0.01$ for $k=2$. The observed robustness against all these errors stems from the independence of Hawking temperature on various details of the black hole.
@@
Finally, we extend our analysis to include two-particle dynamics. Fig. 4(b) depicts the outcome of introducing an interaction term $J_{0}V\sum_{l=1}^{N}n_{l}n_{l+1}$ between two particles. Our results underscore the resilience of our method against attractive interactions, i.e. negative values of $V$, which encourage particle clustering, causing them to act as a single entity, effectively replicating the single-particle behaviour. In contrast, for repulsive interactions ($V=1,2$), the Hawking temperature consistently surpasses its theoretical values across various ranges of $\alpha$.
@@
Conclusions:– Our study utilises the power of locally driving the sites of an optical lattice to generate an efficient analogue quantum simulator for the intricate physics of Dirac particles in a black hole. We introduced a direct approach for determining the Hawking temperature of the simulated black hole from readily accessible onsite atom population measurements at the horizon. The resulting Hawking temperatures exhibit a remarkable resilience against the common sources of errors inherent in Floquet-driven optical lattices, as corroborated by our numerical findings. Additionally, we have unveiled that the presence of weak interatomic interactions introduces a minor, albeit systematic, deviation in the measured value of $T_{H}$.
@@
Exploring black holes beyond the semiclassical limit poses formidable theoretical challenges due to quantum interactions that are intrinsic to quantum gravity. Since optical lattices offer a natural platform for implementing tunable (non-gravitational) interactions [45], our simulator can combine thermalising (scrambling) quantum behaviour with the background curvature of a black hole. Our approach can furthermore be directly extended to higher dimensions, see [51] for an example of a (2+1)D geometry, capitalising on the experimental accessibility of higher-dimensional optical lattices. Finally, combined with quantum gas microscopes, optical lattices naturally provide access to multi-particle correlations, thus offering an ideal system to monitor the behaviour of quantum information during the evaporation of black holes encoded in our simulator by changing the tunnelling coefficients in a time-dependent fashion.
@@
Acknowledgements:– We wish to thank Matthew Horner, F. Nur Ünal, Mehedi Hasan and Dan E. Browne for useful discussions. A.B. acknowledges funding from the EPSRC Centre for Doctoral Training in Delivering Quantum Technologies at UCL, Grant No. EP/S021582/. G.N. acknowledges funding from the Royal Society Te Apārangi and the Cambridge Trust. U.S. acknowledges funding from the EPSRC Programme Grant DesOEQ (Grant No. EP/P009565/1) and the EPSRC Hub in Quantum Computing and Simulation EP/T001062/1. J.K.P. acknowledges support by EPSRC, Grant No. EP/R020612/1.

### v2

@@ II.1 Hawking temperature from radiation population
To demonstrate this, we plot in Fig. 3(c) the time evolution of the population density $n_{h}=\langle c_{j_{h}}^{\dagger}c_{j_{h}}\rangle$ at the event horizon site $j_{h}$. The density $n_{h}$ is initially zero until the wave packet has reached the horizon, after which it reaches a maximum at a time $t_{f}$ before decreasing again. In Fig. 3(d) we plot $t_{f}$ against the temperature $\tilde{T}_{H}$ extracted from the thermalised spectrum and find the simple inverse relationship
@@
with a proportionality constant $\gamma$, directly linking Hawking temperature and time dynamics. This effect also transpires in the inset of Fig. 3(c), where the density evolutions at the event horizon are collapsed by a linear rescaling of the time axis. Note that $\gamma$ depends on the particular experimental conditions, in particular the initial position of the walker. We highlight that monitoring density at a single site is sufficient since the Hawking temperature is controlled by the local curvature at the horizon [3, 4, 6, 5, 88]. Alternatively, one can measure the total probability density outside the horizon, i.e. $P_{\textrm{out}}=\sum_{j=j_{h}+1}^{j=L}n_{j}$, where $n_{j}=\langle c_{j}^{\dagger}c_{j}\rangle$, which provides a direct measure for the total intensity of the Hawking radiation. Within the linear profile of Eq. (3), scaling the encoded Hawking temperature $T_{H}$ amounts to an equivalent global scaling of the tunnelling coefficients $\kappa_{j\>j+1}$, which is directly equivalently to an inverse scaling of time and hence explains the behaviour observed in (8).
@@
So far we have only considered the single-particle case and would hence not yet expect any thermalisation or scrambling. However, due to the specific tunnelling profile that simulates the black-hole horizon, those components of the quantum state that tunnel through the horizon do appear thermal for the degrees of freedom outside the black hole even for a pure initial quantum state. This corresponds to the common notion that the Hawking radiation loses memory of the initial state of the black hole when observed from the outside. This is manifested in our simulator by obtaining a thermal spectrum outside the black hole with the same Hawking temperature regardless of the initial position of the walker inside the black hole.
@@
Figure 4: (a) Error resilience to random on-site potentials (blue) with $\varepsilon=0.1$, additive random nearest-neighbour (nn) tunnelling error (orange) ($\varepsilon=0.1$) and random next-nearest neighbour (nnn) tunnelling with $\varepsilon=0.01$ (green). These results are averaged over time, initialised sites and 100 realisations for a system with size $L=100$ and horizon $j_{h}=50$. The error bar indicates the standard deviation across different realisations. The theoretical Hawking temperature value is represented by the black lines. (b) Density-density interaction of strength $V$ for a system size of $L=30$ and $j_{h}=15$.
III Resilience to errors
@@
We now investigate the sensitivity of the extracted Hawking temperature to errors pertinent to Floquet-driven optical lattice experiments. Specifically, we extract the Hawking temperature $\tilde{T}_{H}$ as per the method used in Fig. 3(b), after introducing noise to the system Hamiltonian in Eq. (1). We model random on-site potentials as an additional term of the form $J_{0}\sum_{j=1}^{N}\varepsilon_{j}\hat{n}_{j}$ in the Hamiltonian with the coefficients $\varepsilon_{j}$ drawn randomly from a standard normal distribution with mean zero and variance $\varepsilon$. Despite the existence of random potentials, the essential features of the extracted $T_{H}$ persist even with an impurity strength of $\varepsilon=0.1$ for a wide range of $\alpha$’s, as shown in Fig. 4(a). We next evaluate nearest-neighbour ($k=1$), and next-nearest neighbour ($k=2$) tunnelling errors, represented by $J_{0}\sum_{j=1}^{N}\varepsilon_{j}(\hat{c}_{j}^{\dagger}\hat{c}_{j+k}+\hat{c}_{j+k}^{\dagger}\hat{c}_{j})$. These errors can reflect higher-order corrections beyond Eq. (6), see [64]. Fig. 4(a) shows that the extracted Hawking temperature remains stable for error strength $\varepsilon=0.1$ for $k=1$ and $\varepsilon=0.01$ for $k=2$. The observed robustness against all these errors stems from the independence of Hawking temperature on various details of the black hole.
@@
Finally, we extend our analysis to include two-particle dynamics. Fig. 4(b) depicts the outcome of introducing an interaction term $J_{0}V\sum_{j=1}^{N}\hat{n}_{j}\hat{n}_{j+1}$ between two particles. Our results underscore the resilience of our method against attractive interactions, i.e. negative values of $V$, which encourage particle clustering, causing them to act as a single entity, effectively replicating the single-particle behaviour. In contrast, for repulsive interactions ($V=1,2$), the Hawking temperature consistently surpasses its theoretical values across various ranges of $\alpha$.
IV 2D Implementation of the black hole simulator
The Floquet simulator can directly be extended to a two-dimensional configuration. This allows probing of non-trivial quantum features that are not present in the one-dimensional case. This includes hallmarks such as the area law for the Bekenstein-Hawking entropy, namely that the black-hole entropy scales with the size of its boundary according to thermodynamic arguments [89, 90, 91, 92, 93]. This property also arises in the Ryu–Takayanagi formula conjectured to describe the relationship between the entanglement entropy of a conformal field theory and the geometry of an Anti-De-Sitter spacetime, i.e. the AdS-CFT correspondence [94, 95, 96, 97, 98]. In combination with the contact interactions readily available with cold atoms, the 2D situation also enables studying the interplay between the locally chaotic, thermalising dynamics of a 2D Hubbard model [99, 100] with the unique global features of the simulated black-hole horizon.
Figure 5: (a) Analogue quantum simulation of a 2D black hole using a Floquet-driven square optical lattice. Sites are driven with amplitudes $A_{n}$ where $n\in\{1,...,17\}$ as labelled and arrows of the same colour indicate the same amplitude. Neighbouring sites are driven $\pi$ out of phase as depicted by opposite arrow directions. The simulated event horizon is at the dashed grey line and the centre of the black hole $(x,y)=0$ is at the site driven by the black arrow with amplitude $A_{1}$. Pictured is the top right quarter of the simulator; the simulation is symmetric under $90\degree$ rotations around the centre. (b) The driving amplitudes indexed by $n$ are chosen such that the tunnelling amplitudes along the $45\degree$ path (blue line in (a)) follow a linear gradient. The alternating sign of $A_{n}$ shown in (b) indicates that neighbouring sites are driven $\pi$ out of phase. (c) The resulting tunnelling coefficients $\kappa_{ij}$ according to Eq. (6) vs. the Euclidean distance $r_{ij}$ of each tunnelling link to the black hole centre. The tunnelling coefficients for the blue line give a perfectly linear profile. Tunnelling coefficients following the $x$ axis (orange line) are also approximately linear. Sites marked by a black circle in (a) have been changed to ensure that all plaquettes have an even number of negative tunnelling links and hence zero flux.
We can simulate a 2D black hole at the origin by approximately encoding a linear curvature profile $f(r)=\alpha(r-r_{h})$, where $r$ denotes the Euclidean distance from the centre of the black hole and the event horizon is at $r_{h}$. The combined index $j=(x_{j},y_{j})$ now corresponds to both $x$ and $y$ coordinates of site $j$. The resulting effective Hamiltonian is then again given by Eq. (1), but with tunnelling along two directions. The renormalised tunnelling coefficients then follow $\kappa_{ij}\propto(r_{ij}-r_{h})$, where $r_{ij}=\sqrt{(x_{i}/2+x_{j}/2)^{2}+(y_{i}/2+y_{j}/2)^{2}}$ is the distance between the tunnelling link connecting sites $i$ and $j$ and the black-hole centre.
As an example, Fig. 5 illustrates an implementation of a 2D black-hole simulator on a square lattice using a Floquet-driven 2D lattice as given by Eq. (4) with appropriately chosen driving amplitudes $A_{j}$. To estimate the required amplitudes, we first fix the tunnelling elements $\kappa_{ij}$ on the blue path along the diagonal in Fig. 5 to match the desired linear gradient $\kappa_{ij}=\alpha(r_{ij}-r_{h})$. For the $N$ sites along the diagonal, Eq. (6) provides a set $\{A_{n}\},n\in\{1,\dots,N\}$ of $N$ corresponding driving amplitudes. Similarly to 1D, neighbouring sites are driven out of phase to each other to minimise the required driving amplitudes, see Figure 5(b). For all remaining sites, we pick that amplitude from the above set $\{A_{n}\}$ that corresponds to the site on the diagonal blue path with the closest matching Euclidean distance to the centre. In order to avoid creating unwanted artificial gauge fields [57], the number of links with negative tunnelling amplitudes on each square plaquette is required to be even, i.e. 0, 2, or 4. As a last step, certain sites just outside of the event horizon (marked by black circles in Figure 5(a)) are hence changed to ensure this condition. Figure 5(c) shows the resulting tunnelling amplitudes for all links as a function of the Euclidean distance from the link to the centre of the black hole, highlighting the approximate linear relationship.
The Hawking temperature $T_{H}$ is encoded in the 2D tunnelling profile analogously to the 1D case, i.e. by the prefactor $\alpha=4\pi T_{H}$. In a square lattice, we do not expect to perfectly preserve the rotational symmetry of the black hole. However, in the limit of large $r_{h}$ or small $d$, the small distortions visible in Figure 5(c) will disappear. An analytical and numerical investigation of similarly encoded linear tunnelling profiles and the robust validity of the resulting encoded Hawking temperature is given in Ref. [62].
V Chaotic behaviour
A central but conjectured property of real black holes is that they are the fastest scrambling—and hence most chaotic—systems in nature [14, 101, 13], a property stemming from the interactions between the particles in the black hole. Crucially, our simulation platform can be extended to interacting quantum particles in the curved spacetime of a black hole by making use of the widely tunable contact interactions natural for fermionic and bosonic cold atoms [58]. This leads to a Hubbard model with an on-site interaction term in addition to the spatially-dependent hopping term from Eq. 1 [102]. Interacting 2D Hubbard models are non-integrable and generally show thermalising, ergodic dynamics [99, 100, 103, 104].
While the experimental implementation can directly use fermionic or bosonic atoms with tunable interactions, we now consider hard-core bosons, i.e. bosonic atoms in the strongly-interacting limit where the interactions suppress site occupations higher than one, for numerical convenience. In this limit, the Bose-Hubbard Hamiltonian is reduced back to the form of Eq. 1, but with the ladder operators $\hat{c}_{i}^{\dagger}$, $\hat{c}_{j}$ now describing hard-core bosons with the commutation relations
$[c_{i},c_{j}^{\dagger}]=[c_{i},c_{j}]=[c_{i}^{\dagger},c_{j}^{\dagger}]=0,\quad\forall i\neq j$ (9)
$\{c_{i},c_{i}^{\dagger}\}=1,\quad(c_{i})^{2}=(c_{i}^{\dagger})^{2}\quad\forall i.$ (10)
This model is expected to be ergodic in the homogeneous case where all tunnellings are equal, corresponding to flat space [99, 103, 104]. To demonstrate the ergodicity of this model also in the black-hole setting, we turn to its level statistics and analyse the ratio between energetically adjacent energy gaps as introduced in Ref. [105]. Namely we calculate
$r_{n}=\frac{\text{min}(g_{n},g_{n-1})}{\text{max}(g_{n},g_{n-1})},$ (11)
where $g_{n}=E_{n+1}-E_{n}\geq 0$ is the gap between two adjacent eigenvalues of the Hamiltonian at a given fixed number of particles, giving $r_{n}\in[0,1]$. In an ergodic system in the thermodynamic limit, the distribution of $r_{n}$ will show level repulsion and follow the Wigner-Dyson statistics characteristic of the Gaussian Orthogonal Ensemble (GOE) of random matrix theory [106, 107, 108, 105]. However, in an integrable system where one would not expect scrambling, there is no level repulsion and the resulting distribution is Poissonian.
Figure 6 compares the mean adjacent energy gap ratios $\langle r\rangle=\frac{1}{s-2}\sum_{n=1}^{s-1}r_{n}$, where $s$ is the Hilbert space dimension, between multiple 2D systems for varying particle number $N$. The orange data corresponds to a system with homogeneous tunnelling, i.e. encoding a flat spacetime, while blue data shows the black hole case with site-dependent tunnellings following Eq. (1). Both the curved and flat systems effectively simulate an $8\times 8$ lattice using cone-like boundary conditions, see [64]. In a finite-sized system, Hamiltonian symmetries cause degeneracies in the spectrum, perturbing the level statistics. To circumvent this, we introduce a local potential of size $\varepsilon$ on a single site of the lattice [64].
The flat system is expected to be ergodic in the thermodynamic limit [103, 104]. The data in Figure 6 shows that our black-hole simulator gives level statistics that are consistently closer to an ergodic distribution compared to the flat system for the same parameter values, thereby demonstrating the potential to study chaotic and scrambling dynamics in this platform.
Figure 6: Comparing level statistics between finite 2D square lattices encoding flat (orange) and curved (blue) space-time backgrounds. The mean gap ratio $\langle r\rangle$ is computed using Eq. 11 for varying particle number $N$. For the blue data, $\alpha=0.46$, $J_{0}=1$ and we add a perturbation of strength $\epsilon=0.01$ (dashed lines) or $\epsilon=0.0625$ (solid lines) to lift degeneracies. For the flat system, homogeneous tunnelling is set to $\kappa J_{0}=0.28$. The expected mean gap ratios $\langle r_{\text{GOE}}\rangle=0.5295\pm 0.0006$ and $\langle r_{\text{Poiss}}\rangle\approx 0.386$ [105] for a GOE and Poissonian ensemble in the thermodynamic limit are indicated by red-dotted and black-dotted lines. The curved spacetime encoded in the blue data gives rise to higher values of $\langle r\rangle$ and hence does not break the ergodicity expected for the flat system.
VI Conclusions
@@
Our study utilises the power of locally driving the sites of an optical lattice to generate an efficient analogue quantum simulator for the intricate physics of Dirac particles in a black hole in one and two spatial dimensions. We introduced a direct approach for determining the Hawking temperature of the simulated black hole from readily accessible onsite atom population measurements at the horizon. The simulated Hawking temperatures are not directly related to the physical temperature of the optical lattice, but can be varied across a broad range by tuning system parameters. This versatility enables the exploration of various scrambling and thermalisation regimes, independent of the limitations imposed by experimental temperature conditions. The resulting Hawking temperatures exhibit a remarkable resilience against the common sources of errors inherent in Floquet-driven optical lattices, including the presence of temperature effects, as corroborated by our numerical findings.
Furthermore, optical lattices offer a natural platform for implementing tuneable (non-gravitational) interactions [58]. Our simulator can therefore probe thermalising (scrambling) quantum behaviour in 2D with the background curvature of a black hole. By analysing level statistics, we demonstrate that adding the background curvature strengthens the ergodicity expected in the flat lattice. Together with the realisation of Hawking radiation, being able to implement such chaotic behaviour could enable us to realise the Hayden-Preskill quantum teleportation protocol responsible for the information retrieval from evaporating black holes [109, 110, 111].
@@
Finally, combined with quantum gas microscopes, optical lattices naturally provide access to multi-particle correlations, thus offering an ideal system to monitor the behaviour of quantum information during the evaporation of black holes encoded in our simulator by changing the tunnelling coefficients in a time-dependent fashion.
Acknowledgements.
@@
We wish to thank Matthew Horner, F. Nur Ünal, Mehedi Hasan and Dan E. Browne for useful discussions. A.B. acknowledges funding from the EPSRC Centre for Doctoral Training in Delivering Quantum Technologies at UCL, Grant No. EP/S021582/. G.N. acknowledges funding from the Royal Society Te Apārangi and the Cambridge Trust. U.S. acknowledges funding from the EPSRC Programme Grant DesOEQ (Grant No. EP/P009565/1) and the EPSRC Hub in Quantum Computing and Simulation EP/T001062/1. J.K.P. acknowledges support by EPSRC, Grant No. EP/R020612/1.